Solve -2z - 43 = -17 using inverse operations.: Solve -2z - 43 = -17 using inverse operations.
Answer by oscargut(666)

2z - 43 = -17
2z-43+43=-17+43
2z=26
(1/2)2z=(1/2)26
z=13